# Attention Meets Post-hoc Interpretability: A Mathematical Perspective

### v1

@@ 1 Introduction
In Section 1.1, we discuss the relevant literature, in particular focusing on the debate around attention-based explanations. We describe the model that we study in Section 2. In Section 3 we specifically discuss attention-based explanations. In Section 4 (resp. Section 5), we derive expressions for the gradient (resp. LIME) explanations associated to our model. These expressions (Theorems 4.1 and 5.1) are explicit with respect to the model parameters and the input document, thus allowing us to pinpoint exactly the differences between these approaches. We draw our conclusion in Section 6. All our theoretical claims are supported by mathematical proofs and empirical validation, detailed in the Appendix. The code for the model and the experiments are available at https://github.com/gianluigilopardo/attention_meets_xai.
@@ 1.1 Related work
It is important to note that there is currently no definitive theoretical support for either side of the debate on whether attention serves as an explanation. The positions presented by both Jain & Wallace (2019) and Wiegreffe & Pinter (2019) are primarily based on empirical experiments. The subsequent debate provide valuable insights and provoke thoughtful discussion, but do not conclusively prove or disprove the interpretability of attention mechanisms.
@@ 2.2 Attention
Figure 3: Attention matrices across the heads. Each head is represented by a distinct matrix, demonstrating the unique focus each head has on different parts of the document. The matrices illustrate that tokens within the document can carry significantly different weights, indicating the varying importance or relevance of each token in the context of the document. The aggregation of these weights to provide token-level weights is a critical aspect.
@@
Limitations.
The main differences between our model and practical architectures are the following: (i) number of layers, (ii) skip connections, (iii) non-linearities. (i): we only consider a single layer, which already brings non-linearity while giving a quite challenging analysis and a very good performance in practice for our task. (ii): we do not consider skip connections since we did not observe an increase in performance while adding them, but our analysis can easily be adapted to this setting since this operation is linear. (iii): we refrain from incorporating additional non-linearities, i.e., we do not introduce a ReLU layer (as suggested by Vaswani et al. (2017)) or any of its variants.
@@ 4.2 Gradient of our model
[…]
@@ 5.2 Limit explanations for our model
Figure 4: Illustration of the accuracy of Eq. (19). Each token corresponds to a unique word of the document. The boxplots correspond to $5$ runs of LIME with default parameters, whereas the red crosses correspond to the predictions given by Theorem 5.1. The document $\xi$ has $T=d=99$ tokens and is classified as negative.
@@
Theorem 5.1 is proved in Section B. The challenging part of the proof is to derive good approximations for Eq. (17), since the model we consider, although single-layered, is highly non-linear. We nevertheless observe a very good match between our theoretical approximation and the empirical outputs of LIME with default parameters, which we illustrate in Figure 4 on a particular example, and in a more quantitative way in Appendix B.
@@
We now make a few comments. (i) it is clear from Eq. (19) that LIME explanations are quite different from gradient-based explanations (Eq. (14)), with the exception of the leading term (which is proportional to $\alpha_{t}W_{\ell}v_{t}$) which we recognize in both expressions. (ii) one can see that LIME explanations are approximately an affine transformation of the $\alpha_{t}^{(i)}$. (iii) as it is the case for gradient-based explanation, there is a major difference with plain attention-based explanations: the last layer comes into account in the explanation. To put it plainly, let us assume that $K=1$ and that $W_{\ell}v_{t}=0$, the influence of $\alpha_{t}$ disappears whatever its value. We see this as an advantage for LIME, since this situation corresponds to the model killing the influence of token $t$ in later stages, although a positive attention score is given. (iv) from Eq. (19), we see that LIME explanations will be near zero whenever $\alpha_{t}^{(i)}v_{t}^{(i)}\approx\alpha_{h,t}^{(i)}v_{h,t}^{(i)}$, a scenario in which the attention $\times$ value of head $i$ given to token $t$ if comparable to that of the attention given to the [UNK] token. This makes a lot of sense, while calling for a careful choice of embedding for the replacement token.
@@
We conclude this section by discussing limitations of Theorem 5.1. First, as in previous work, the approximation is only true for large document and window size. This comes without surprise, but we note that the results are experimentally satisfying for documents which are a few dozen tokens long. Second, we assume in Theorem 5.1 that all tokens are distinct. We conjecture that this assumption can be relaxed if the maximal multiplicity of words belonging to the local dictionary is not too large. Experimentally, Eq. (19) holds for documents containing repeated words which is why we formulate in that way, allowing for repetitions.
@@ 6 Conclusion and future work
In this paper, we offered a theoretical analysis on how post-hoc explanations relates to a single-layer multi-head attention-based network. Our work contributes to the ongoing debate in this area by providing exact and approximate expressions for post-hoc explanations. Through these expressions, we were able to highlight the fundamental differences between attention, gradient, and perturbation-based explanations.
It is crucial to acknowledge that the quest for perfect explanations remains elusive; no single method has emerged as entirely satisfactory. However, it is clear that current models employ attention scores in a non-intuitive manner to arrive at the final prediction. In particular, these scores go through a series of further transformations, which is ignored when looking solely at attention scores. These scores also always provide a positive explanation, in contrast to perturbation-based and gradient-based approaches. For these reasons, we believe that they can extract more valuable insights than a mere examination of attention weights. This finding aligns with the assertions made by Bastings & Filippova (2020).
As future work, we plan to expand our analysis to other post-hoc methods, such as Anchors. We also would like to obtain similar statements (connecting explanations to the parameters of the model) for more complicated architectures, including skip connections, additional non-linearities, and multi-layer models. Finally, there is some interplay between the sampling mechanism of perturbation-based methods (often replacing at the word level) and the tokenizer used by the model (tokens are often subwords) which we would like to understand better.
@@
We empirically validate the accuracy of Theorem 5.1 by computing the norm-$2$ error between the LIME weights from the official implementation (available on Github at https://github.com/marcotcr/lime) and our approximation. The norm-$2$ error, computed over the full test set (see Section E), is $0.808$, with a standard deviation of $0.219$.
@@ D.2 Probability computations
Appendix E Experiments

### v2

@@ 1 Introduction
In Section 1.1, we discuss the relevant literature, in particular focusing on the debate around attention-based explanations. We describe the model that we study in Section 2. In Section 3 we specifically discuss attention-based explanations. In Section 4 (resp. Section 5), we derive expressions for the gradient (resp. LIME) explanations associated to our model. These expressions (Theorems 4.1 and 5.1) are explicit with respect to the model parameters and the input document, thus allowing us to pinpoint exactly the differences between these approaches. In Section 6, we discuss the main limitations of our work, including the theoretical assumptions underlying the model under examination. We draw our conclusion in Section 7. All our theoretical claims are supported by mathematical proofs and empirical validation, detailed in the Appendix. The code for the model and the experiments are available at https://github.com/gianluigilopardo/attention_meets_xai.
@@ 1.1 Related work
It is important to note that there is currently no definitive theoretical support for either side of the debate on whether attention serves as an explanation. The positions presented by both Jain & Wallace (2019) and Wiegreffe & Pinter (2019) are primarily based on empirical experiments. The subsequent debate provided valuable insights and provoked thoughtful discussion, but did not conclusively prove or disprove the interpretability of attention mechanisms.
[…]
@@ 2.2 Attention
Figure 3: Attention matrices across the heads. Each head is represented by a distinct matrix, demonstrating the unique focus each head has on different parts of the document. The matrices illustrate that tokens within the document can carry significantly different weights, indicating the varying importance or relevance of each token in the context of the document. The aggregation of these weights to provide token-level scores is a critical aspect. Note that Eqs. (10) and (11) correspond to the average and the maximum values, respectively, of the first row across all six matrices.
@@
We discuss limitations and the main differences between our model and practical architectures in Section 6.
@@ 4.2 Gradient of our model
$\left.+\frac{\alpha_{t}^{(i)}}{\sqrt{d_{\text{att}}}}W_{\ell}^{(i)}\left(v_{t}^{(i)}-\sum_{s=1}^{T_{\max}}\alpha_{s}^{(i)}v_{s}^{(i)}\right)(W_{k}^{(i)})^{\top}q\right]\in\mathbb{R}^{d_{e}}\,.$
@@ 5.2 Limit explanations for our model
Figure 4: Illustration of the accuracy of Eq. (5.1). The boxplots show the results from $5$ runs of LIME with default parameters, while the red crosses indicate the predictions given by Theorem 5.1. The document $\xi$ contains $T=99$ tokens and $d=71$ distinct words and is classified as a negative review. Note that Theorem 5.1 holds true even with $T\neq d$ for reasonable word multiplicities, as discussed in Section 6.
@@
Theorem 5.1 is proved in Section B. The challenging part of the proof is to derive good approximations for Eq. (17), since the model we consider, although single-layered, is highly non-linear. Note that Theorem 5.1 relies on the assumption that all tokens are distinct. While we conjecture that this assumption can be relaxed (as discussed in Section 6), it is necessary for a rigorous comparison between attention and LIME explanations. However, our experiments were conducted without assuming all tokens are distinct. We nevertheless observe a very good match between our theoretical approximation and the empirical outputs of LIME with default parameters, which we illustrate in Figure 4 on a particular example, and in a more quantitative way in Appendix B.
@@
We now make a few comments. (i) it is clear from Eq. (5.1) that LIME explanations are quite different from gradient-based explanations (Eq. (14)), with the exception of the leading term (which is proportional to $\alpha_{t}W_{\ell}v_{t}$) which we recognize in both expressions. (ii) one can see that LIME explanations are approximately an affine transformation of the $\alpha_{t}^{(i)}$. (iii) as it is the case for gradient-based explanation, there is a major difference with plain attention-based explanations: the last layer comes into account in the explanation. To put it plainly, let us assume that $K=1$ and that $W_{\ell}v_{t}=0$, the influence of $\alpha_{t}$ disappears whatever its value. We see this as an advantage for LIME, since this situation corresponds to the model killing the influence of token $t$ in later stages, although a positive attention score is given. (iv) from Eq. (5.1), we see that LIME explanations will be near zero whenever $\alpha_{t}^{(i)}v_{t}^{(i)}\approx\alpha_{h,t}^{(i)}v_{h,t}^{(i)}$, a scenario in which the attention $\times$ value of head $i$ given to token $t$ if comparable to that of the attention given to the [UNK] token. This makes a lot of sense, while calling for a careful choice of embedding for the replacement token.
6 Limitation
The main differences between the model described in Section 2 and practical architectures are the following: (i) number of layers, (ii) skip connections, (iii) non-linearities. (i): we only consider a single layer, which already brings non-linearity while giving a quite challenging analysis and a very good performance in practice for our task. (ii): we do not consider skip connections since we did not observe an increase in performance while adding them, but our analysis can easily be adapted to this setting since this operation is linear. (iii): in line with typical theoretical works (Gunasekar et al., 2018), we refrain from incorporating additional non-linearities; specifically, we do not introduce a ReLU activation or any of its variants.
The primary limitation of the present analysis lies in its focus on a single-layer model. This facilitates a deep theoretical exploration, yet its applicability to more intricate architectures may not be direct. The extension of this analysis to a multi-layer architecture introduces additional theoretical challenges. In the context of multi-layer transformers, approximation errors have the potential to accumulate and intensify as they traverse through the network. Moreover, the assumptions applicable to a single layer may not necessarily hold true for deeper networks, especially those incorporating non-linearities such as ReLU activations. However, even for simple architectures, many questions remain unresolved, and the interpretability of these models has not been thoroughly studied in a formal manner.
Several theoretical studies on transformers share these limitations. For example, Jelassi et al. (2022) elucidates how Vision Transformers discern spatial patterns within a single-layer, single-head architecture. Similarly, Tarzanagh et al. (2023) establishes the correspondence between the optimization geometry of self-attention and an SVM problem. Von Oswald et al. (2023) suggests that training Transformers with a specific objective can induce a form of meta-learning, exemplified on a linear single-layer model. Additionally, Edelman et al. (2022) investigate transformers in time series forecasting, showcasing their superiority over traditional methods by accurately capturing temporal dependencies. In a bid to enhance transformer efficiency and scalability, Fu et al. (2023) introduce sparse attention mechanisms and efficient training strategies by formalizing single-layer transformers. Furthermore, Makkuva et al. (2024) tackle transformer interpretability, developing tools to visualize and comprehend attention patterns within the models, aiming to bridge the gap between high performance and the imperative for transparency in critical applications. In Appendix E, we report experiments on a multi-layer transformer.
@@ 6 Limitation
We conclude this section by discussing limitations of Theorem 5.1. First, as in previous work, the approximation is only true for large document and window size. This comes without surprise, but we note that the results are experimentally satisfying for documents which are a few dozen tokens long. Second, we assume in Theorem 5.1 that all tokens are distinct. This assumption, though a simplification, allows for a rigorous formalization of LIME’s behavior using its default parameters as defined in the official implementation. Experimentally, Eq. (5.1) holds for documents containing repeated words. We validated Theorem 5.1 disregarding this assumption, by computing the norm-$2$ error between the official LIME weights and our approximation (as illustrated in Figure 4): the average error over the full test set (described in Appendix F) is $0.808$, with a standard deviation of $0.219$. From a theoretical aspect, we conjecture that this assumption can be relaxed if the maximal multiplicity of tokens is small relative to $T$. If many tokens are identical, this is no longer true: consider, for instance, the extreme case of two groups of identical tokens. We delve into this topic in more detail in Appendix B.1.
@@ 7 Conclusion and future work
In this paper, we offered a theoretical analysis on how post-hoc explanations relates to a single-layer multi-head attention-based network. Our work contributes to the ongoing debate in this area by providing exact and approximate expressions for post-hoc explanations on such model. Through these expressions, we were able to highlight the fundamental differences between attention-based, gradient-based, and perturbation-based explanations. This deeper understanding not only enriches the ongoing discourse surrounding interpretability but also offers valuable insights for practitioners and researchers navigating the complexities of transformers’ interpretation.
It is crucial to acknowledge that the quest for perfect explanations remains elusive; no single method has emerged as entirely satisfactory. However, it is clear that current models employ attention scores in a non-intuitive manner to arrive at the final prediction. In particular, these scores go through a series of further transformations, which is ignored when looking solely at attention scores. These scores also always provide a positive explanation, in contrast to (most) perturbation-based and gradient-based approaches. For these reasons, we believe that they can extract more valuable insights than a mere examination of attention weights. This finding aligns with the assertions made by Bastings & Filippova (2020).
As future work, we plan to broaden the scope of our analysis by extending our investigations to diverse range of post-hoc interpretability methods, including Anchors, thus understanding model explanations across different methodologies. We also would like to obtain similar statements (connecting explanations to the parameters of the model) for more complicated architectures, including skip connections, additional non-linearities, and multi-layer models, enabling us to discern the relationship between model parameters and different explanations. Addionally, there is some interplay between the sampling mechanism of perturbation-based methods (often replacing at the word level) and the tokenizer used by the model (tokens are often subwords) which we would like to understand better. Lastly, we emphasize that our focus in this paper has been on text classification. This choice allows us to capitalize on well-established, and broadly studied post-hoc explainers and conduct a thorough theoretical analysis based on this specific domain. However, we intend to expand the scope of applications for our analysis. Specifically, we remark that our study focused on token-level explanations. Moving forward, we intend to extend our findings beyond text models to encompass other domains, such as computer vision.
Acknowledgements
This work has been supported by the French government, through the NIM-ML project (ANR-21-CE23-0005- 01), and by EU Horizon 2020 project AI4Media (contract no. 951911). Most of this work was realized while DG was employed at Université Côte d’Azur.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Acknowledgements
B.1 Discussion on Theorem 5.1
A theoretical limitation of Theorem 5.1 it the assumption of distinct tokens. This assumption, while technically a simplification, enables a rigorous formalization of LIME’s behavior using its default parameters as defined in the official implementation. We conducted quantitative experiments that disregarded this assumption, and the results still hold. We empirically validate the accuracy of Theorem 5.1 by computing the norm-$2$ error between the LIME weights from the official implementation (available on Github at https://github.com/marcotcr/lime) and our approximation. The average norm-$2$ error, computed over the full test set (see Section F), is $0.808$, with a standard deviation of $0.219$.
The primary challenge in proving a formal result that allows for repetitions lies in the use of Lemma D.1. The key intuition in the current proof mechanism is that if only one element of the denominator of $A_{t}$ varies randomly, this has a minimal overall effect on the entire denominator, given that it has $T\gg 1$ terms. However, if many tokens are identical, this is no longer true, and it can result in high variance (consider, for instance, the extreme case of two groups of identical tokens), which prevents us from using Lemma D.1. Nevertheless, we conjecture that if the tokens are not distinct, but the maximal multiplicity of tokens is small relative to $T$, our findings hold true (and this is empirically true).
On a more practical note, we highlight that by default, LIME-text perturbs input data by removing all occurrences of individual words or characters (bow=True, i.e., bag-of-words), and this is the subject of our study. However, if the underlying model uses word location (as in our classifier), a possibility is to set bow=False (as recommended in their notebook), so that any occurrence of the same word is considered a distinct token. By using this option, the same applies in Theorem 5.1: the same words in different parts of the text are considered different tokens.
@@ D.2 Probability computations
Appendix E Experiments on multi-layer architecture
Figure 6: Relation between LIME weights and attention for a $6$-layer $6$-head attention-based classifier. Attention weights correspond to the average attention over the $6$ heads of the first layer. Error bars represent the standard deviation of LIME weights on $5$ repetition.
We have conducted numerical experiments on a multi-head, multi-layer architecture. We trained a classifier with $6$ layers and $6$ heads on the IMDb dataset (refer to Appendix F), achieving an accuracy of $82.22\%$. Our interest lies in exploring the relationship between LIME and the attention weights. We measured the correlation between LIME coefficients and the $\alpha$-avg (refer to Eq. (5.1)) for the first layer. An illustration is available at Figure 6, where the document corresponds to Figure 4 of our paper. The Pearson’s correlation in this case is $\rho=-0.424$. We attribute the negative sign to the document being classified as negative (as in Figure 4). Attention weights consistently fall within the range of $[0,1]$. Considering the absolute values, the rankings of the two explanations are closely aligned, and the correlation is $\rho=0.672$. Although we cannot explicitly state the dependency for multi-layers as in Eq. (5.1), our experiments suggest a significant relationship. We conclude that the attention weights are interconnected with LIME coefficients, which adapt more effectively to the model. We are currently conducting broader experiments and will incorporate their results and subsequent discussions into the manuscript.
Appendix F Experiments